⦁ What is the molarity of a solution that contains 10.0 grams of Silver Nitrate that has been dissolved in 750g of water


1
Expert's answer
2022-03-03T05:42:01-0500

1. C = (m / M) / (V / 1000) = (10 / 170) / (750 / 1000) = 0.078 M